Jane is x years old today. Her brother Kenny is 4 years older. After seven years, their combined age will be 24 years.
Write a linear equation for their total combined age after 7 years.
Find Janes age today.

Answer by candydreams99(2)

Well, if Jane is x years old that means Kenny is x+4.
Seven years later, Jane would be x+7 and Kenny would be x+11.
It says their combined age would be 24, so the equation would be:
x+7 + x+11 = 24
2x + 18=24
(subtract 18 from both sides)
- 18 + 2x + 18=24 +-18
2x = 6
(divide each side by 1/2 to get x by itself)
x = 3
x+4 = 7
So, Jane is 3 and Kenny is 7.